I already submitted this question. The missing data is Apha=.05. Please help me. I am totally lost and need your help. Thank you again

1. The EPA is concerned about the quality of drinking water served on airline flights. A sample of 158 flights yielded poor quality water on 20 flights. At alpha=.05, does this indicate that more than 10% of flights have poor-quality water?
2. According to the sample, 12.66% (20/158) of the flights served poor-quality water. Can you simply report to the EPA that more than 10% of all airline flights serve poor-quality water? Why or why not?
3. If an error occurred, what type of error was it?

Answer by stanbon(48568)

1. The EPA is concerned about the quality of drinking water served on airline flights.
A sample of 158 flights yielded poor quality water on 20 flights. At alpha=.05, does this indicate that more than 10% of flights have poor-quality water?
--------
Ho: p = 0.10
Ha: p > 0.10 (right-tail test with alpha = 5%)
---------------
Critical value = 1.645
--------------------
p-hat = 20/158 = 0.1266
Test Statistic:
z(0.1266) = (0.1266-0.10)/sqrt(0.1*0.9/158)= 46.69777...
--------------------------
Since the test statistic is greater than the critical value
reject Ho. There is strong statistical evidence that the
proportion pf poor quality is not 10%
2. According to the sample, 12.66% (20/158) of the flights served poor-quality water. Can you simply report to the EPA that more than 10% of all airline flights serve poor-quality water? Why or why not?
Since this is a test of Ho that p = 10% you can only report that the
test shows significant statistical evidence that the proportion is not 10%.
---------------------------
3. If an error occurred, what type of error was it?
Type I; rejecting Ho when in fact it is true.
=======================
Cheers,
Stan H.